# Hessian spectrum at the global minimum of the spherical pure-like mixed $p$-spin glasses

### v1

@@ 3 Pure-like Mixture models
$\mathbb{E}\operatorname{Crt}_{N}(B,D)$
$=C_{\xi,N}\int_{B\times D}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
@@
Firstly, we present the main result of this section as below.
Theorem 3.1 (Pure-like mixture models).
where $y_{0}$ is defined in (3.7).
@@
Note that the function $\Psi_{*}$ is Lipschitz and denoted by $L_{\Psi_{*}}$ the Lipschitz constant. A little algebra yields that the claim (3.13) holds for sufficiently small $\varepsilon$ with
@@
By the key observation (3.13) in Lemma 3.2, for sufficiently small $\varepsilon$, it follows that $\lim_{N\rightarrow\infty}I_{N}^{1}=0$, which together with (3.21) finishes the lemma. ∎
Finally, we finish the proof of Theorem 3.1 with the help of Lemmas 3.3 and 3.4.
@@
Note that $GS=-E_{0}$ as given in (3.8) for the $pure$-$like$ mixture models under Assumption I. The proof of this theorem follows a similar argument as in Theorem 2.1. We only need to replace the parameters $c_{p}$ by $y_{0}$ and $r_{p}$ by $2\sqrt{\xi^{\prime\prime}}$, since we can use Lemmas 3.3 and 3.4 to play the role of Lemmas 2.2 and 2.3 in Theorem 2.1, respectively. ∎
@@ Acknowledgements.
Hao Xu would like to thank his previous supervisor Qiang Zeng for introducing us to consider the problem of Hessian spectrum at the global minimum of spherical mixed $p$-spin models. His careful reading and numerous suggestions have greatly improved the quality of this paper.

### v2

@@ 3 Pure-like Mixture models
$\mathbb{E}\operatorname{Crt}_{N}(B,D)$ $=C_{\xi,N}\int_{B\times D}\exp\left\{-\frac{N}{2}(x,y)\Sigma^{-1}(x,y)^{\top}\right\}$
@@
Note that the function $\Psi_{*}$ is Lipschitz and denote by $L_{\Psi_{*}}$ the Lipschitz constant. A little algebra yields that the claim (3.11) holds for sufficiently small $\varepsilon$ with
@@
By the key observation (3.11) in Lemma 3.1, for sufficiently small $\varepsilon$, it follows that $\lim_{N\rightarrow\infty}I_{N}^{1}=0$, which together with (3.19) finishes the proof of this lemma. ∎
Finally, we prove Theorem 1.2 with the help of Lemmas 3.2 and 3.3.
@@
Note that $GS=-E_{0}$ as given in (3.8) for the $pure$-$like$ mixture models under Assumption I. The proof of this theorem follows similar arguments as those in Theorem 1.1. We only need to replace the parameters $c_{p}$ by $y_{0}$ and $r_{p}$ by $2\sqrt{\xi^{\prime\prime}}$, since Lemmas 3.2 and 3.3 can serve the same role in Theorem 1.2 as Lemmas 2.1 and 2.2 do in Theorem 1.1, respectively. ∎
@@ Acknowledgements.
The authors would like to thank Qiang Zeng for introducing them to the problem of Hessian spectrum at the global minimum of mixed spherical $p$-spin models. His careful reading and numerous suggestions have greatly improved the quality of this paper.